Angular displacement is 2t^2-9t +4 radius find the angular velocity and angular acceleration
1
Expert's answer
2021-08-10T08:01:42-0400

By defintion, the angular velocity and angular acceleration are the first and second derivative of angular displacement "\varphi(t) = 2t^2 - 9t + 4" with respect to time.

Thus, obtain velocity:


"\omega (t) = (2t^2 - 9t + 4)' = 4t - 9"

and acceleration:


"\varepsilon(t) = (2t^2 - 9t + 4)'' = (4t - 9)' = 4"